# Understanding Long Videos with Multimodal Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have allowed recent LLM-based approaches to achieve excellent performance on long-video understanding benchmarks. We investigate how extensive world knowledge and strong reasoning skills of underlying LLMs influence this strong performance. Surprisingly, we discover that LLM-based approaches can yield surprisingly good accuracy on long-video tasks with limited video information, sometimes even with no video specific information. Building on this, we exploring injecting video-specific information into an LLM-based framework. We utilize off-the-shelf vision tools to extract three object-centric information modalities from videos and then leverage natural language as a medium for fusing this information. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across multiple video understanding benchmarks. Strong performance also on robotics domain tasks establish its strong generality. Our code will be released publicly.
@@ 1 Introduction
In this work, we systematically study this question in the context of video question-answering (QnA) benchmarks, building two modality-constrained baselines to highlight our findings. These two frameworks are tagged Just-LLM and Single-Frame-VLM. The first is constrained to access only the task textual query allowing it to answer using only world knowledge (i.e. no task-specific visual information). The latter is given access to task context with an additional single center-frame from the video as input. We discover how these models perform significantly better than random prediction on multiple long-video understanding benchmarks (see Table 1). In fact, the latter utilizing purely world knowledge and contextual information even outperforms multiple recent state-of-the-art video understanding works (see Table 2), challenging the notion of how much video information is actually utilized by existing approaches to solve these complex video QnA tasks.
We next focus on efficient inference to allow rapid experimentation with our LLM based frameworks. Therein, we explore suitable prompting and templating to adapt likelihood selection techniques from prior work (Robinson et al., 2023) to video QnA tasks. Our resulting framework achieves more efficient inference with competing performance in comparison to prior work that commonly use auto-regressive generation to tackle long-video QnA benchmarks (Zhang et al., 2023a; Balavzevi’c et al., 2024).
Motivated by our initial findings on modality-constrained performance, we study how to inject additional video-specific information into our framework using natural language in an interpretable manner in order to further improve video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information using off-the-shelf vision tools, and utilize natural language to fuse this multi-modal information using templating operations. Our resulting approach, termed Multi-Modal Video Understanding (MVU) framework, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation. Moreover, our proposed MVU exhibits generality with its strong performance even on robotics domain tasks.
@@
    Uncover surprisingly strong performance on complex video-language tasks by modality-constrained baselines with no access to video-specific information.
@@
    Propose a novel VLM-based video QnA framework that extracts video specific object-centric information followed by natural language based fusion.
Extensive evaluation of our framework on video QnA benchmarks as well as robotics domain tasks establish its clear benefit. We also highlight how our proposed MVU is a framework (as opposed to a fixed model) that can easily be integrated with any newer, stronger LLMs or VLMs to utilize their improved capabilities.
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just-LLM utilizes only world knowledge with zero task-specific awareness. (left-right) Single-Frame-VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
@@ 2 Related Work
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Early works explore querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019) followed by focus on temporal modeling of sequential events (Lei et al., 2018; 2020; Hosseini et al., 2022; Xiao et al., 2022a; b). While motivated by these works, MVU integrates such object information with large language models (LLMs) in a zero-shot manner requiring no video-level training. More recent works leverage LLMs (Yu et al., 2023b; Papalampidi et al., 2023; Wang et al., 2022; Balavzevi’c et al., 2024; Yang et al., 2022; Ye et al., 2023b) to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Limited zero-shot operation is explored in Wang et al. (2023); Zhang et al. (2023a), but we differ in using object-centric modality information, efficient LLM sampling, and zero video level training.
Large Language Model Reasoning: Recent LLMs (OpenAI, 2023; Chowdhery et al., 2022; Chiang et al., 2023) demonstrate multiple forms of strong reasoning abilities (Kıcıman et al., 2023; Creswell & Shanahan, 2022; Liu et al., 2023b) including combining different information (Weston & Sukhbaatar, 2023). Their recent open-source variants (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2023) achieve equally promising skills using scaled-down models (Jiang et al., 2023) while also demonstrating strong world knowledge (Yu et al., 2023a; AlKhamissi et al., 2024; Zhao et al., 2023; Wang & Zhao, 2023; Xu et al., 2024; Li et al., 2023d) even in domains such as robotics Li et al. (2024). In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their abilities for video QnA tasks.
@@ 3.2 Likelihood Selection
In light of this, we explore prompting and likelihood calculation techniques optimal for applying Likelihood Selection on long video QnA tasks with either LLMs or even VLMs. Adapting this technique unlocks autoregressive LLMs / VLMs ability to solve multiple selection problems with only one forward pass. This is in contrast to next token sampling requiring iterative generations dependent on previous outputs for each answer token. As illustrated in Figure 2, likelihood selection combines each question $x_{t}$ with all $M=5$ candidate answers $y_{1,...,5}$ to calculate the conditional likelihood $P(y_{i}|x_{t})$ for each answer. This is calculated in a single LLM forward pass and selects the answer with the highest conditional likelihood among all the answers. We direct the reader to complete details of our video QnA specific prompting templates and calculations in Table A.4.
In addition to the considerable inference speed-up from likelihood selection, we also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks (see Hanu et al. (2023)). We empirically validate the improved performance from such behaviour in our ablations (see Tab. 6.).
@@ 3.3 Modality Constrained Variants
We hypothesize that Just-LLM with no access to task-specific knowledge is constrained to generate predictions utilizing its internal world knowledge (e.g. physics, culture, human common sense). We refer to this as world modality. For a given question regarding a natural video and a set of candidate answers, there is a possibility that one choice is more probable given how our world operates. In cases that this choice turns out to be correct, the internal world knowledge of the LLM allows it to easily select that choice resulting in above random performance. This variant of our framework highlights such cases in long video QnA tasks.
@@
We evaluate these two modality-constrained variants and summarize our findings in Table 1. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. In the case of Just-LLM variant, we achieve performance significantly higher than random selection (+25.8% on EgoSchema-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QnA benchmarks that can be answered correctly purely using world knowledge. We also highlight our Single-Frame-VLM performing on par with state-of-the-art LLM based approach from Zhang et al. (2023a). In particular, for EgoSchema-S we outperform Zhang et al. (2023a) which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve, possibly in an interpretable manner.
Table 1: Modality Constrained Variants: We evaluate on ES-S (public subset of EgoSchema (Mangalam et al., 2023)) and NextQA-T (test set of NextQA (Xiao et al., 2021)) datasets. We report accuracy (%) and inference time per sample (s). Recent state-of-the-art approach from Zhang et al. (2023a) using multi-frame inputs is also reported for comparison.
Method Param ES-S NextQA-T
@@
Random - 20.0 - 20.0 -
Just-LLM 7B 45.8 0.41 40.1 0.55
SF-VLM 13B 55.8 1.89 51.2 2.03
SOTA 20B 50.8 381 54.3 207
@@ 4.2 Object-Centric Information Modalities
This pipeline for extracting per-frame information using an image-trained VLM closely resembles prior work such as (Zhang et al., 2023a). While motivated by such work, we explore the direction of how more fine-grained information could be extracted from videos to solve these tasks more efficiently. Given the role of object interactions in defining the various actions and events in videos, we hypothesize that extracting object-centric information followed by explicit capturing of their temporal dependencies (as opposed to generic frame-level descriptions) would be highly beneficial to solving video tasks.
We also highlight how the use of vision tools to extract specific forms of information provides an additional level of interpretability to the functions of our overall pipeline. We next discuss how we fuse each of these information modalities using natural language as a medium.
@@ 4.3 Language based Fusion
where ${x}_{v}^{c}$ is the center frame extracted from the video $x_{v}$. We refer the reader to Appendix A for further details on prompting and template operations used as well as examples for our language-based fusion.
5 Experiments
In this section, we first discuss our experimental setup and datasets. Next, we present our findings when operating in modality-constrained settings, followed by evaluations of our approach on multiple video question-answering benchmarks to demonstrate its performance. Finally, we perform ablations of our framework to verify the usefulness of all proposed components.
5.1 Long Video Question Answering
Table 2: Ego-Schema Dataset Evaluation: We report top-1 accuracy (%) for video question answering on the open-subset (500 videos) and closed full-set (8564 videos) of Ego-Schema (Mangalam et al., 2023). Our proposed MVU achieves state-of-the-art performance on this benchmark under zero-shot operation with no video level training. We also draw attention to our modality-constrained baselines, Vid-LLM and Vid-VLM, that achieve surprisingly competitive performance on this benchmark. Approaches utilizing proprietary LLMs with significantly higher parameters (e.g. GPT-4) are de-emphasized.
@@ 5.1 Long Video Question Answering
Params Frames Subset Full
Random Selection - - - 0 20.0 20.0
ImageViT (Papalampidi et al., 2023) 1B 16 40.8 30.9
ShortViViT (Papalampidi et al., 2023) 1B 16 47.9
SeViLA+ShortViViT (Papalampidi et al., 2023) 5B 32 49.6 31.3
LongViViT (Papalampidi et al., 2023) 1B 256 56.8 33.3
MC-ViT-L (Balavzevi’c et al., 2024) 424M 128 62.6 44.4
FrozenBiLM (Yang et al., 2022) 1.2B 90 - 26.9
VIOLET (Fu et al., 2022) 198M 5 - 19.9
mPLUG-Owl (Ye et al., 2023b) 7.2B 5 - 31.1
InternVideo (Wang et al., 2022) 478M 90 32.1
SeViLA (Yu et al., 2024) 4B 32 25.7 22.7
LLoVi (Zhang et al., 2023a) 13B 180 50.8 33.5
Vamos (Wang et al., 2023) 13B - - 36.7
LLoVi (Zhang et al., 2023a) 1.8T 180 58.3 50.3
Vamos (Wang et al., 2023) 1.8T - - 48.3
Just-LLM (ours) 7B 0 45.8 29.1
Single-Frame-VLM (ours) 13B 1 55.8 36.4
MVU (ours) 13B 16 60.3 37.6
@@
On EgoSchema dataset, results reported in Table 2 demonstrate the state-of-the-art performance of our framework in comparison to work operating under similar settings. We reiterate how our approach is both zero-shot and requires no video-level training. In contrast, multiple prior zero-shot works utilize pseudo-labels on the downstream task videos for transductive training (Fu et al., 2022; Ye et al., 2023b; Yu et al., 2024). We also highlight how our MVU even outperforms multiple recent supervised (video-caption) approaches (Papalampidi et al., 2023). Among zero-shot methods, while our results are subpar to the trillion parameter variants of (Zhang et al., 2023a; Wang et al., 2023) that utilize proprietary closed language models, we outperform with a clear margin against the fairer comparison to their similar parameter variants (more in Appendix C).
Table 3: Next-QA Dataset Evaluation: We report top-1 accuracy (%) for the test-split of Next-QA dataset (Xiao et al., 2021). Our proposed MVU achieves state-of-the-art results under zero-shot settings with no video-level training. We also highlight our competitive performance across the dataset’s causal, why, and how sub-categories. In table header, ZS stands for zero-shot and VT stands for video level training.
Method ZS VT Params Causal Why How All
@@
MVU (ours) 13B 55.7 48.2 64.2 55.4
We next evaluate our framework on the NextQA benchmark and report these results in Table 3. Our MVU achieves competitive results under zero-shot settings. While Yu et al. (2024) outperforms our approach, we note how they require video-caption localization pretraining and appears to overfit to this dataset considering their relatively lower performance on other datasets (see Table 2).
While these two datasets focus on MCQ style QnA, we also explore the generality of our MVU framework on open-ended style QnA tasks. We detail these evaluations in Appendix G, highlighting how MVU performance extends beyond MCQ style tasks.

### v3

@@ Abstract
Large Language Models (LLMs) have allowed recent LLM-based approaches to achieve excellent performance on long-video understanding benchmarks. We investigate how extensive world knowledge and strong reasoning skills of underlying LLMs influence this strong performance. Surprisingly, we discover that LLM-based approaches can yield surprisingly good accuracy on long-video tasks with limited video information, sometimes even with no video specific information. Building on this, we exploring injecting video-specific information into an LLM-based framework. We utilize off-the-shelf vision tools to extract three object-centric information modalities from videos and then leverage natural language as a medium for fusing this information. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across multiple video understanding benchmarks. Strong performance also on robotics domain tasks establish its strong generality. Code: github.com/kahnchana/mvu
@@ 1 Introduction
In this work, we systematically study this question in the context of video question-answering (QnA) benchmarks, building two modality-constrained baselines to highlight our findings. These two frameworks are tagged Just-LLM and Single-Frame-VLM. The first is constrained to access only the task textual query (i.e. no task-specific visual information). The latter is given access to task context with an additional single center-frame from the video as input. We discover how these models perform significantly better than random prediction on multiple long-video understanding benchmarks (see Table 1, similar findings in Min et al. (2024)). In fact, the latter utilizing purely world knowledge and contextual information even outperforms multiple recent state-of-the-art video understanding works (see Table 2), challenging the notion of how much video information is actually utilized by existing approaches to solve these complex video QnA tasks.
We next focus on efficient inference to allow rapid experimentation with our LLM based frameworks. Therein, we explore suitable prompting and templating to adapt likelihood selection techniques from prior work (Robinson et al., 2023) to video QnA tasks. Our resulting framework achieves more efficient inference with improved performance in comparison to prior work that commonly use auto-regressive generation to tackle long-video QnA benchmarks (Zhang et al., 2023a; Balavzevi’c et al., 2024; Wang et al., 2025).
Motivated by our initial findings on modality-constrained performance, we study how to inject additional video-specific information into our framework using natural language in a concise and interpretable manner to further improve video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information using off-the-shelf vision tools, and utilize natural language to fuse this multi-modal information using templating operations. Our resulting approach, termed Multi-Modal Video Understanding (MVU) framework, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation. Moreover, our proposed MVU exhibits generality with its strong performance even on robotics domain tasks.
@@
    Uncover surprisingly strong performance on complex video-language tasks by modality-constrained baselines with limited access to video-specific information.
@@
    Novel VLM-based video QnA framework that extracts concise video specific object-centric information followed by natural language based fusion.
We integrate our MVU framework over multiple different baselines and obtain performance improvements across 20 different datasets establishing both its effectiveness and generality. Our evaluations are performed zero-shot with no video-level training on these datasets which cover video QnA tasks (short, medium, and long videos) as well robotics domain tasks.
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just-LLM utilizes only world knowledge with zero task-specific awareness. (left-bottom) Single-Frame-VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
@@ 2 Related Work
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Early works explore querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019) followed by focus on temporal modeling of sequential events (Lei et al., 2018; 2020; Hosseini et al., 2022; Xiao et al., 2022a; b). While motivated by these works, MVU integrates such object information with large language models (LLMs) in a zero-shot manner requiring no video-level training. More recent works leverage LLMs (Yu et al., 2023b; Papalampidi et al., 2023; Wang et al., 2024b; Balavzevi’c et al., 2024; Wang et al., 2024a) to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Zero-shot operation is explored in Wang et al. (2023a); Zhang et al. (2023a); Min et al. (2024); Wang et al. (2023b; 2024c), but we differ in using object-centric modality information and efficient LLM sampling.
Large Language Model Reasoning: Recent LLMs (OpenAI, 2023; Chowdhery et al., 2022; Chiang et al., 2023) demonstrate multiple forms of strong reasoning abilities (Kıcıman et al., 2023; Creswell & Shanahan, 2022; Liu et al., 2023b) including combining different information (Weston & Sukhbaatar, 2023). Their recent open-source variants (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2023) achieve equally promising skills using scaled-down models (Jiang et al., 2023) while also demonstrating strong world knowledge (Yu et al., 2023a; AlKhamissi et al., 2024; Zhao et al., 2023; Wang & Zhao, 2023; Xu et al., 2024; Li et al., 2023d) even in domains such as robotics (Li et al., 2024). In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their abilities for video QnA tasks.
@@ 3.2 Likelihood Selection
In light of this, we explore prompting and likelihood calculation techniques optimal for applying Likelihood Selection on long video QnA tasks with either LLMs or even VLMs. Adapting this technique unlocks autoregressive LLMs / VLMs ability to solve multiple selection problems with only one forward pass as illustrated in Figure 2. This is in contrast to next token sampling requiring iterative generations dependent on previous outputs for each answer token. This process uses a likelihood measure based on the LLM latent space allowing better semantic awareness compared to exact or template matching. In addition to the candidate answer batching, we follow prior work to include all candidates in the prompt as well. We direct the reader to Table A.4 for complete details on semantic awareness, candidates in prompts, and video QnA specific implementation.
In addition to the considerable inference speed-up from likelihood selection, we also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks (see Hanu et al. (2023)). We empirically validate the improved performance from such behavior in our ablations (see Table 7).
@@ 3.3 Modality Constrained Variants
We hypothesize that Just-LLM with no access to task-specific knowledge is constrained to generate predictions utilizing its internal world knowledge (e.g. physics, culture, human common sense). We refer to this as world modality. For a given question regarding a natural video and a set of candidate answers, there is a possibility that one choice is more probable given how our world operates. In cases that this choice turns out to be correct, the internal world knowledge of the LLM allows it to easily select that choice resulting in above random performance. This variant of our framework highlights such cases in long video QnA tasks. A similar baseline is used in Min et al. (2024).
@@
We evaluate these two modality-constrained variants and summarize our findings in Table 1. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. In the case of Just-LLM variant, we achieve performance significantly higher than random selection (+25.8% on ES-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QnA benchmarks that can be answered correctly purely using world knowledge. We also highlight our Single-Frame-VLM performing on par with state-of-the-art LLM based approach from Zhang et al. (2023a). In particular, for ES-S we outperform Zhang et al. (2023a) which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve, possibly in an interpretable manner.
Table 1: Modality Constrained Variants: We report accuracy (%) and inference time per sample (s) on the public subset of EgoSchema (ES-S) and test set of NextQA (NextQA-T) datasets. Note that recent state-of-the-art from Zhang et al. (2023a) (SOTA) and our variants are implemented with common LLMs / VLMs and evaluated under identical settings.
Method Param Video Frames ES-S NextQA-T
@@
Random - - 20.0 - 20.0 -
Just-LLM 7B 0 45.8 0.41 40.1 0.55
SF-VLM 13B 1 55.8 1.89 51.2 2.03
SOTA 20B 180 50.8 381 54.3 207
@@ 4.2 Object-Centric Information Modalities
This pipeline for extracting per-frame information using an image-trained VLM closely resembles prior work such as (Zhang et al., 2023a). While motivated by such work, we explore the direction of how more fine-grained information could be extracted from videos to solve these tasks more efficiently. Given the role of object interactions in defining the various actions and events in videos, we hypothesize that extracting object-centric information (as opposed to generic frame-level descriptions) followed by modeling of their temporal dependencies would provide more concise representations better suited to efficiently solve these tasks.
@@ 4.3 Language based Fusion
where ${x}_{v}^{c}$ is the center frame extracted from the video $x_{v}$ (more details in Appendix A). In comparison to prior work such as Zhang et al. (2023a), we note that our fused information is more concise allowing better utilization of the fixed context length in an LLM (see Appendix K for more details).
Table 2: Ego-Schema Dataset Evaluation: We report top-1 accuracy (%) for video question answering on Ego-Schema (Mangalam et al., 2023) test set (5031 videos). Our proposed MVU achieves state-of-the-art performance on this benchmark under zero-shot operation with no video level training. We also draw attention to our modality-constrained SF-VLM baseline that achieves surprisingly competitive performance.
@@
Closed
Model
Params Full
Random Selection - - - - 20.0
VIOLET (Fu et al., 2022) 198M 19.9
FrozenBiLM (Yang et al., 2022) 1.2B 26.9
SeViLA (Yu et al., 2024) 4B 22.7
mPLUG-Owl (Ye et al., 2023b) 7.2B 31.1
InternVideo (Wang et al., 2022) 478M 32.1
ImageViT (Papalampidi et al., 2023) 1B 30.9
SeViLA+ShortViViT (Papalampidi et al., 2023) 5B 31.3
LongViViT (Papalampidi et al., 2023) 1B 33.3
MC-ViT-L (Balavzevi’c et al., 2024) 424M 44.4
InternVideo2 (Wang et al., 2024b) 7B 55.8
Tarsier (Wang et al., 2024a) 7B 49.9
Tarsier (Wang et al., 2024a) 34B 61.7
Vamos (Wang et al., 2023a) 13B 36.7
LLoVi (Zhang et al., 2023a) 13B 33.5
LangRepo (Kahatapitiya et al., 2024) 12B 41.2
Vamos (Wang et al., 2023a) 1.8T 48.3
LLoVi (Zhang et al., 2023a) 1.8T 50.3
LifelongMemory (Wang et al., 2023b) 1.8T 62.4
MoreVQA (Min et al., 2024) - 51.7
VideoAgent (Wang et al., 2025) 1.8T 54.1
VideoTree (Wang et al., 2024c) 1.8T 61.1
LVNet (Park et al., 2024) 1.8T 61.1
SF-VLM (ours) 13B 36.4
SF-VLM + MVU (ours) 13B 37.6
LVNet + MVU (ours) 1.8T 61.3
5 Experiments
In this section, we first discuss our experimental setup and datasets. Next, we evaluate MVU on multiple video question-answering and robotics task benchmarks followed by ablative studies.
5.1 Long Video Question Answering
@@
On EgoSchema dataset, results reported in Table 2 demonstrate the state-of-the-art performance of our framework. We integrate MVU over SF-VLM and LVNet (Park et al., 2024) baselines for fair comparison to work operating under different settings. We reiterate how our approach is both zero-shot and requires no video-level training (and our selected baselines are similar). In comparison to prior work utilizing open models, our SF-VLM+MVU achieves clear performance improvements, even out-performing works using video-caption supervision for training (Papalampidi et al., 2023; Balavzevi’c et al., 2024). Compared to methods utilizing proprietary closed language models extending to trillion parameter scale (Zhang et al., 2023a; Wang et al., 2023a; Min et al., 2024; Wang et al., 2025), our LVNet+MVU variant using similar scale achieves improved performance. We also implement several such large-scale approaches under scaled-down common settings as our smaller variant (details in Appendix C). Here also we achieve clear performance gains.
Table 3: Next-QA Dataset Evaluation: We report top-1 accuracy (%) for the Next-QA dataset (Xiao et al., 2021). Our proposed MVU achieves state-of-the-art results under zero-shot settings with no video-level training. In table header, ZS stands for zero-shot and VT stands for video level training.
Method ZS VT Params Cau. Tem. Des. All
@@
ViperGPT (Surís et al., 2023) 175B - - - 60.0
LLoVi (Zhang et al., 2023a) (GPT-4) 1.8T 69.5 61.0 75.6 67.7
MoreVQA (Min et al., 2024) 1.7T 70.2 64.6 - 69.2
VideoAgent (Wang et al., 2025) 1.7T 72.7 64.5 81.1 71.3
VideoTree (Wang et al., 2024c) 1.7T 75.2 67.0 81.3 73.5
LVNet (Park et al., 2024) 1.8T 75.0 65.5 81.5 72.9
SF-VLM + MVU (ours) 13B 55.7 48.2 64.2 55.4
LVNet + MVU (ours) 1.8T 75.2 66.8 81.3 73.3
We next evaluate our framework on the NextQA benchmark and report these results in Table 3. We similarly integrate MVU with two baselines. Our MVU achieves state-of-the-art results under zero-shot settings. While Yu et al. (2024) outperforms our approach, we note how they require video-caption localization pretraining and appears to overfit to this dataset considering their relatively lower performance on other datasets (see Table 2).
We also evaluate MVU on the LongVideoBench dataset which contains even longer videos and present these results in Appendix H. While these three datasets focus on MCQ style QnA, we also explore the generality of our MVU framework on open-ended style QnA tasks in Appendix G.